Simplify the rational expression. If the rational expression cannot be simplified, so state.

?2 − 25

?−5

I am not sure of the correct steps to solve this problem. I got x+5 but I am not sure that is correct. Can you break down the problem, step by step.
Answer #1

Solved; Since CQ²-62) = (att) (a-6) .: X2 - 25 = x2 – 5² - (Octs) (C-5) Answes.